OpenStudy (anonymous):

Evaluate the expression log(base3)9+2log(base3)3

OpenStudy (mertsj):

\[\log_{3}9+2\log_{3}3=2+2(1)=4 \]

OpenStudy (anonymous):

Thanks a bunch! =)